Hello, I hope you can help me solve this. Please forgive me if I don't write it in the appropriate manner:
What is the inverse of 2 raised to the power of 8 divided by 2 raised to the power of 2?
Thank you for your help.
Regards
Lisa McCallum

Answer by stanbon(48535)

What is the inverse of 2 raised to the power of 8 divided by 2 raised to the power of 2?
---------
2^8/2^2 = 2^(8-2)=2^6
The inverse would be 1/2^6 or 2^-6
Cheers,
Stan H.